OpenStudy (compassionate):

Completing squares. Lets says I have x^2 - 4x = 0 To complete the square, I know I do something with -4x. Is it take the square root of 4?

OpenStudy (cggurumanjunath):

(x-(4/2))^2=0+(4/2)^2

OpenStudy (cggurumanjunath):

(x-2)^2=4

hartnn (hartnn):

You need to divide the co-efficient of x by 2. then square it. Add it on both sides.

OpenStudy (aravindg):

You add a term every time. It is is given an (coefficient of x term/2)^2.

OpenStudy (compassionate):

-4/2 = -2 Then the square of -2 is -1... Isn't that an imaginary number?

OpenStudy (aravindg):

Squiare of -2 is 4 o.o

OpenStudy (compassionate):

You can't take roots of negative numbers....

hartnn (hartnn):

square of -2 is -2 times -2

OpenStudy (compassionate):

Oh...

OpenStudy (compassionate):

So... (x^2 - 4x) + (y^2 + 8y) = -4 I would say... (x^2 - 4x + 4) + (y^2 + 8y + 16) = - 4 + 4 + 16...

OpenStudy (aravindg):

Yep.

OpenStudy (compassionate):

Then I would say... (x - 2)^2 + (y + 4)^2 = 16...

hartnn (hartnn):

=4^2 yeah, thats an equation of circle with centre 2,-4 and radius = 4

OpenStudy (aravindg):

Yes!

OpenStudy (compassionate):

Danke schone. Mucho gracias. شكرا.